The table shows the percentage P of an employee's annual salary S, in thousands of dollars, spent on the process of recruiting and training a replacement.

S = annual salary
(thousands of dollars)
P = percentage of
salary for recruiting
20 15
30 16
40 19
60 45
70 50

Find the equation of the regression line. (Round regression line parameters to two decimal places.)

P(S) =

What cost does the regression line give for replacement of an employee who makes an annual salary of $50,000? (Round your answer to the nearest dollar.)
Answer #1

Here S (salary) is the independent variable and P (percentage) is the dependent variable.

We know that, trend line = P = a + b*S

Where,

a = {(∑P)*(∑S2) – (∑S)*(∑S*P)} / {(n)*(∑S2) – (∑S)*(∑S)}

b = {(n)*(∑S*P) – (∑S)*(∑P)} / {(n)*(∑S2) – (∑S)*(∑S)}

S (IN 1000s)

P

S*P

S^2

20

15

300

400

30

16

480

900

40

19

760

1600

60

45

2700

3600

70

50

3500

4900

220

145

7740

11400

Here,

∑P = 145

∑S = 220

∑S*P = 7740

∑S2 = 11400

n = 5

Putting everything to find a and b

a = {(145*11400) – (220*7740)}/{(5*11400) – (220*220)}

= (1653000 – 1702800)/(57000 – 48400)

= -49800/8600

= -5.79

a = -5.79

b = {(5*7740) – (220*145)}/{(5*11400) – (220*220)}

= (38700 – 31900)/(57000 – 48400)

= 6800/8600

= 0.79

b = 0.79

Hence, P = 0.79*S – 5.79

For S = 50 (in 1000s)

P = 0.79*50 – 5.79 = 39.5 – 5.79 = 33.71 = 34

Hence For Salary 50000, P = 34

IF YOU LIKE THE ANSWER, PLEASE GIVE AN UP-VOTE OR THUMB UP. THIS WILL ENCOURAGE ME TO ANSWER MORE!!